# Regularization of the ensemble Kalman filter using a non-parametric, non-stationary spatial model

### v3

@@ Abstract
The sample covariance matrix of a random vector is a good estimate of the true covariance matrix if the sample size is much larger than the length of the vector. In high-dimensional problems, this condition is never met. As a result, in high dimensions the EnKF ensemble does not contain enough information to specify the prior covariance matrix accurately. This necessitates the need for regularization of the analysis (observation update) problem. We propose a regularization technique based on a new spatial model on the sphere. The model is a constrained version of the general Gaussian process convolution model. The constraints on the location-dependent convolution kernel include local isotropy, positive definiteness as a function of distance, and smoothness as a function of location. The model allows for a rigorous definition of the local spectrum, which, in addition, is required to be a smooth function of spatial wavenumber. We regularize the ensemble Kalman filter by postulating that its prior covariances obey this model. The model is estimated online in a two-stage procedure. First, ensemble perturbations are bandpass filtered in several wavenumber bands to extract aggregated local spatial spectra. Second, a neural network recovers the local spectra from sample variances of the filtered fields. We show that with the growing ensemble size, the estimator is capable of extracting increasingly detailed spatially non-stationary structures. In simulation experiments, the new technique led to substantially better EnKF performance than several existing techniques.
@@
Process convolution model, Spatial non-stationarity, Data assimilation, Ensemble Kalman Filter, Neural Bayes estimator
@@ 1 Introduction
Modern data assimilation in geosciences and related areas relies on the ensemble approach, in which the probability distribution of the unknown system state is represented by a finite sample (ensemble) of pseudo-random realizations (ensemble members). This allows for realistic, spatially varying prior covariances in the analysis (the observation update step of the data assimilation cycle). In practical applications, the most widely-used approach is the Ensemble Kalman Filter (EnKF) and its numerous modifications and extensions111 See A for a more technical description of ensemble filtering in the linear setting (which is sufficient for our purposes in this study). . The principal problem of the ensemble approach is that maintaining a large ensemble in the course of data assimilation (sequential filtering) is computationally expensive. For real-world systems modeled with high-resolution models with billions of degrees of freedom, this means that only relatively small ensembles are affordable. As a result, such ensembles can provide the analysis with only scarce information on the prior distribution. The sample covariance matrix is a poor estimate of the true covariance matrix if the sample size is much lower than the dimensionality of the state space. This requires a kind of regularization, that is, the introduction of additional information on the prior covariances. The following techniques have been proposed so far.
@@
    The most popular approach is covariance localization (tapering), (e.g., Furrer and Bengtsson, 2007, Houtekamer and Zhang, 2016), which reduces spurious long-distance correlations through element-wise multiplication of the sample covariance matrix by an ad-hoc positive definite localization matrix. This technique efficiently removes a lot of noise in the sample covariance matrix. However, it cannot cope with the noise at small distances (where it is the strongest). And it inevitably leads to an underestimation of the assumed length scale of the forecast-error field and thus to imbalances between different fields (Lorenc, 2003).
@@
    Blending sample covariances with static (time-mean) covariances is widely used in meteorological ensemble-variational schemes (Buehner et al., 2013, Lorenc et al., 2014). In statistical literature, similar techniques are known as shrinkage estimators (Ledoit and Wolf, 2004). Sample covariances are noisy but they do contain the useful flow-dependent “signal”. Static covariances are noise-free but can be irrelevant under some flow conditions and observation coverage. A convex combination of the two covariance matrices proved to be useful (see the above references) but it is not selective: the noise in the sample covariances is reduced to the same extent as the flow-dependent non-stationary signal.
@@
    A similar technique is temporal averaging of sample covariances (i.e., blending with recent past covariances). Berre et al. (2015) and Bonavita et al. (2016) used prior ensembles from several previous assimilation cycles to increase the ensemble size. Lorenc (2017) used time-shifted ensemble perturbations for the same purpose. Tsyrulnikov and Rakitko (2017) arrived at this technique by assuming that the true covariance matrix is an unknown random matrix and introducing a secondary filter in which the covariances are updated.
@@
    (Ueno and Tsuchiya, 2009) proposed to regularize the sample covariance matrix by imposing a sparse structure in the inverse covariance (precision) matrix. A similar approach was taken by Hou et al. (2021). Gryvill and Tjelmeland (2023) imposed a Markov random field structure to obtain a sparse prior precision matrix. Boyles and Katzfuss (2021) postulated sparsity in inverse Cholesky factors of the prior covariance matrix.
@@
    One more option is to adopt a parametric forecast-error covariance model and estimate parameters of the model from the forecast ensemble. Skauvold and Eidsvik (2019) used maximum likelihood estimates of a few model parameters in stationary and non-stationary settings.
@@
In this paper we consider systems whose state vector represents a spatial field, e.g., atmosphere, ocean, and the like. Taking advantage of this representation of the state vector, we propose to regularize the analysis problem by imposing a spatial model. In this regard, our approach is similar to the two latter techniques (items 6 and 7 above). In contrast to parametric covariance models, we aim at a non-parametric formulation to obtain a more flexible model. By non-parametric, we mean defined by parameters whose number is not fixed (as with a parametric model) but grows with the growing spatial resolution of the model. Compared to the wavelet-diagonal models, which provide regularizing information by imposing zero cross-correlations between wavelet coefficients, our model does so by imposing several constraints on the general process convolution model. The two principal constraints are local stationarity (under which a spatially non-stationary process can be approximated by a stationary process locally in space) and smoothness of local spectra.
In section 2 we introduce a non-parametric locally stationary Gaussian process convolution model on the sphere (and on the circle). In section 3 we present a computationally efficient estimator of the model from an ensemble of random field realizations. Consistency of the estimator is proven in I. Section 4 describes how the spatial model is used to regularize the analysis step of the EnKF. We test the new approach in numerical experiments with a non-stationary random field model on the sphere (section 5) and in cyclic filtering on the circle (section 6).
@@ 2 Spatial model
As noted in the Introduction, we employ a non-parametric spatial model for the forecast-error random field to regularize the data assimilation problem. Following (Higdon et al., 1999), we rely on the Gaussian process convolution model. In contrast to most applications, which postulate a parametric model for the spatial convolution kernel (e.g., Lemos and Sansó, 2009, Sham Bhat et al., 2012, Heaton et al., 2014, Li and Zhu, 2020), we choose a non-parametric approach to allow for variable shapes of spatial covariances. Non-parametric convolution kernels were considered by Tobar et al. (2015) and Bruinsma et al. (2022) in the stationary time series context. Here we propose a non-parametric non-stationary spatial convolution model on compact manifolds such as the sphere and the circle.
The default domain is the two-dimensional unit sphere ${\mathbb{S}}^{2}$ (the domain of interest in geosciences). We also formulate and test the model on the unit circle ${\mathbb{S}}^{1}$ (used as a much simpler analog of the spherical case). We refer to ${\mathbb{S}}^{2}$ and ${\mathbb{S}}^{1}$ just as the sphere and the circle in the sequel. We denote points on the sphere as $x,y,\dots$ and area elements simply by ${\rm d}x,{\rm d}y,\dots$ On the sphere, we use the terms stationarity and isotropy interchangeably. The terms field and process are also used interchangeably. We confine ourselves to band-limited functions because in the context of sequential filtering, the spatial field to be estimated inevitably has limited spatial resolution. We also assume that probability distributions we deal with in this study have densities with respect to the Lebesgue measure.
@@ 2.1 General process convolution model
The (non-stationary in general) covariance function of the process $\xi(x)$ defined by Eq. (1) is $B(x,x^{\prime})=\operatorname{\mathbb{E}}\xi(x)\xi(x^{\prime})=\int w(x,y)w(x^{\prime},y)\,{\rm d}y.$ Technically, this equation implies that for ${\mathsf{Var}}\,\xi(x)$ to be finite, the kernel $w(x,y)$ needs to be square integrable with respect to its second argument for any $x$: $\int w^{2}(x,y)\,{\rm d}y<\infty$.
@@ 2.2 Constraining the model: strategy
In the formulation Eq. (1), the model is not identifiable, that is, the function $w(x,y)$ is not unique given the process covariance function $B(x,x^{\prime})$. Indeed, consider an isotropic convolution kernel $\psi(\rho(y,y^{\prime}))$ (where $\rho(y,y^{\prime})$ stands for the great-circle distance between the points $y$ and $y^{\prime}$) whose Fourier transform (on ${\mathbb{S}}^{1}$) or Fourier-Legendre transform (on ${\mathbb{S}}^{2}$) is equal to one in modulus (see B for the definition of the Fourier-Legendre transform we use). Then, it is straightforward to see that the covariance function remains unchanged if $w(x,y)$ is convolved with $\psi$. (The convolution with such a kernel is analogous to the multiplication by an orthogonal matrix in the space discrete case.)
@@ 2.3 Locally isotropic kernel
$\xi(x)=\int u(x,\rho(x,y))\,\alpha(y)\,{\rm d}y.$ (3)
In the remainder of this subsection, we develop spectral representations of the process defined by Eq. (3) and its covariance function. We perform the spectral (Fourier-Legendre) expansion (see Eq. (34)) of $u(x,\rho)$ with $x$ being fixed:
@@
where $\sigma_{\ell}(x)$ are real-valued functions (we call them spectral functions or spectral standardd deviations) and $P_{\ell}$ are the Legendre polynomials. Substituting $\rho=\rho(x,y)$ into Eq. (4) and applying the addition theorem of spherical harmonics (B), we obtain
@@
Then, we make use of the spectral expansion of the white noise, Eq. (35), and substitute it along with Eq. (5) into Eq. (3). Utilizing orthonormality of spherical harmonics, we obtain the basic spectral representation of the random process that satisfies Eq. (3):
@@
The spatial covariances can readily be obtained from Eq. (6) by taking into account that all $\widetilde{\alpha}_{\ell m}$ are mutually uncorrelated and again applying the addition theorem of spherical harmonics:
$B(x,x^{\prime})=\operatorname{\mathbb{E}}\xi(x)\xi(x^{\prime})=\sum^{\ell_{\rm max}}_{\ell=0}\frac{2\ell+1}{4\pi}\,\sigma_{\ell}(x)\,\sigma_{\ell}(x^{\prime})\,P_{\ell}(\cos\rho(x,x^{\prime})).$ (7)
If $u(x,\rho)=u(\rho)$ or, equivalently, if all spectral functions $\sigma_{\ell}(x)$ do not depend on $x$, then the model defined by Eq. (3) or Eq. (6) becomes a stationary (isotropic) random field model. Equation (7) shows that the process variance ${\mathsf{Var}}\,\xi(x)$ can be obtained by summing $\frac{2\ell+1}{4\pi}f_{\ell}(x)$ known in the stationary case as the variance spectrum (energy per degree $\ell$). Again in the stationary case, $f_{\ell}$ is often called the modal spectrum (energy per ‘mode”, i.e., per pair $(\ell,m)$). In the non-stationary case, in which $f_{\ell}$ depends on $x$, we call $f_{\ell}(x)=\sigma_{\ell}^{2}(x)$ the local modal spectrum or just the local spectrum.
Non-negative definiteness (i.e., validity) of the model’s non-stationary covariance function $B(x,x^{\prime})$ is established in C.
The formulation of the spatial model Eq. (6) on the circle is outlined in D.

### v5

# ©2024. This manuscript version is made available under the CC-BY-NC-ND 4.0 license https://creativecommons.org/licenses/by-nc-nd/4.0/

@@ Abstract
The sample covariance matrix of a random vector is a good estimate of the true covariance matrix if the sample size is much larger than the length of the vector. In high-dimensional problems, this condition is never met. As a result, in high dimensions the Ensemble Kalman Filter’s (EnKF) ensemble does not contain enough information to specify the prior covariance matrix accurately. This necessitates the need for regularization of the analysis (observation update) problem. We propose a regularization technique based on a new spatial model. The model is a constrained version of the general Gaussian process convolution model. The constraints include local stationarity and smoothness of local spectra. We regularize EnKF by postulating that its prior covariances obey the spatial model. Placing a hyperprior distribution on the model parameters and using the likelihood of the prior ensemble data allows for an optimized use of both the ensemble and the hyperprior. A linear version of the respective estimator is shown to be consistent. A more accurate nonlinear neural-Bayes implementation of the estimator is developed. In simulation experiments, the new technique led to substantially better EnKF performance than several existing techniques.
@@
Process convolution model , Spatial non-stationarity , Data assimilation , Ensemble Kalman Filter , Neural Bayes estimator , Gaussian process
@@
[…]
[main]organization=HydroMeteorological Research Center of Russia, country=Russian Federation
[…]
[scnd]organization=National Research University Higher School of Economics, country=Russian Federation
@@ 1 Introduction
Modern data assimilation in geosciences and related areas relies on the ensemble approach, in which the probability distribution of the unknown system state is represented by a finite sample (ensemble) of pseudo-random realizations (ensemble members). This allows for realistic, spatially varying prior covariances in the analysis (the observation update step of the data assimilation cycle). In practical applications, the most widely-used approach is the Ensemble Kalman Filter (EnKF) and its numerous modifications and extensions222 See section 2 for a brief introduction and A for a more technical description of sequential ensemble filtering in the linear setting (which is sufficient for our purposes in this study). . The principal problem of the ensemble approach is that maintaining a large ensemble in the course of data assimilation (sequential filtering) is computationally expensive. For real-world systems modeled with high-resolution models with billions of degrees of freedom, this means that only relatively small ensembles are affordable. As a result, such ensembles can provide the analysis with only scarce information on the prior distribution. The sample covariance matrix is a poor estimate of the true covariance matrix if the sample size is much lower than the dimensionality of the state space. This requires a kind of regularization, that is, the introduction of additional information on the prior covariances.
1.1 Related work
The following EnKF regularization techniques have been proposed so far.
@@ 1.1 Related work
    The most popular approach is covariance localization (tapering), (e.g. Furrer and Bengtsson, 2007, Houtekamer and Zhang, 2016), which reduces spurious long-distance correlations through element-wise multiplication of the sample covariance matrix by an ad-hoc positive definite localization matrix. This technique efficiently removes a lot of noise in the sample covariance matrix. However, it cannot cope with the noise at small distances (where the noise is the strongest). And it inevitably leads to an underestimation of the assumed length scale of the forecast-error field and thus to imbalances between different fields (Lorenc, 2003).
@@
    Blending prior sample covariances with static (time-mean) covariances is widely used in meteorological ensemble-variational schemes (Buehner et al., 2013, Lorenc et al., 2014). In statistical literature, similar techniques are known as shrinkage estimators (Ledoit and Wolf, 2004). Sample covariances are noisy but they do contain the useful flow-dependent ‘signal’. Static covariances are noise-free but can be irrelevant under some flow conditions and observation coverage. A convex combination of the two covariance matrices proved to be useful (see the above references) but it is not selective: the noise in the sample covariances is reduced to the same extent as the flow-dependent non-stationary signal.
@@
    A similar technique is temporal averaging of prior sample covariances (i.e., blending with recent past covariances). Berre et al. (2015) and Bonavita et al. (2016) used prior ensembles from several previous assimilation cycles to increase the ensemble size. Lorenc (2017) used time-shifted ensemble perturbations for the same purpose. Tsyrulnikov and Rakitko (2017) arrived at this technique by assuming that the true covariance matrix is an unknown random matrix and introducing a secondary filter in which the covariances are updated.
@@
    Ueno and Tsuchiya (2009) proposed to regularize the sample covariance matrix by imposing a sparse structure in the inverse covariance (precision) matrix. A similar approach was taken by Hou et al. (2021). Gryvill and Tjelmeland (2023) imposed a Markov random field structure to obtain a sparse prior precision matrix. Boyles and Katzfuss (2021) postulated sparsity in inverse Cholesky factors of the prior covariance matrix.
@@
    One more option is to adopt a parametric model for the prior covariances and estimate parameters of the model from the forecast ensemble. Skauvold and Eidsvik (2019) used maximum likelihood estimates of a few model parameters in stationary and non-stationary settings.
@@
1.2 Motivation and contributions
In this paper we consider systems whose state vector represents a spatial field, e.g., atmosphere, ocean, and the like. Taking advantage of this interpretation of the state vector, we propose to regularize the analysis problem by imposing a spatial model. In this regard, our approach is similar to the two latter techniques (items 6 and 7 above). In contrast to parametric covariance models, we aim at a non-parametric formulation to obtain a more flexible model. By non-parametric, we mean defined by parameters whose number is not fixed (as with a parametric model) but grows with the increasing spatial resolution of the model. Compared to the wavelet-diagonal models, which provide regularizing information by imposing zero cross-correlations between wavelet coefficients, our model does so by imposing several constraints on the general process convolution model. The two principal constraints are local stationarity (under which a spatially non-stationary process can be approximated by a stationary process locally in space) and smoothness of what we call local spectra.
We assume that the prior distribution of the state at each filtering time step is multivariate Gaussian with the mean equal to the ensemble mean (or, for a linear forecast model, to the deterministic forecast, see Remark in A.4). The prior covariances are postulated to be governed by the spatial model introduced in this paper. We place a hyperprior distribution on the model parameters and use the likelihood of the ensemble data to online estimate the spatial model. The estimated model is then employed to compute a square root of the prior covariance matrix, which is finally used in the analysis.
The reason why we propose to rely on a Bayesian estimate of a spatial model in the EnKF analysis is that a more principled (than the above ad-hoc techniques) way of covariance regularization can lead to a better estimate of the true prior covariance matrix and thus better performance of the filter. It is also worth noting that the introduction of the hyperprior distribution implies an enhancement of EnKF that goes beyond blending of sample covariances with time-mean covariances (mentioned in item 2 above).
Our key contributions are:
  • A constrained non-parametric, non-stationary process convolution model on the sphere. The concept of local stationarity of the resulting spatial process defined by conditioning on the random convolution kernel.
  • An estimator of the spatial model suitable for high-dimensional applications. A proof that a linear version of the estimator is consistent. A nonlinear, neural Bayes realization of the estimator.
  • Implementation of the model and its estimator in EnKF, leading to a new Locally Stationary Ensemble Filter (LSEF).
  • Numerical experiments with simulated data with the LSEF analysis on the sphere and the LSEF on the circle.
The rest of the paper is organized as follows. Section 2 gives a brief overview of sequential ensemble filtering. The theoretical part of the paper consists of two big sections, 3 and 4, supported by several Appendices. In section 3 we introduce a non-parametric locally stationary Gaussian process convolution model on the sphere (and on the circle). In section 4 we propose an estimator of the model from an ensemble of random field realizations. Section 5 describes how the spatial model is used to regularize the analysis step of the EnKF. The experimental part of the paper is presented in sections 6and 7.
@@ 1.2 Motivation and contributions
2 Background on sequential ensemble filtering
The goal of sequential filtering is the estimation of the current state of an evolving system, having (i) a forecast model capable of predicting the future state of the system given its current state, (ii) a set of observations on the system, and (iii) an observation model that relates the observations to the system state (the truth).
In the time-discrete filtering, observations are available at discrete time instants. The estimates of the (ground) truth are sought at the same time instants. A sequential filter consists of alternating short-term forecast (time update) and analysis (observation update) steps. In the time-discrete linear filtering, which is sufficient to test the covariance regularization technique we propose in this research, the evolution of the truth ${\bf x}_{t}$ between consecutive time instants $t$ is assumed to be governed by the stochastic equation
${\bf x}_{t}={\bf F}_{t}{\bf x}_{t-1}+\boldsymbol{\eta}_{t},$ (1)
where ${\bf F}_{t}$ is the known forecast-model operator (a matrix) and $\boldsymbol{\eta}_{t}$ is the unknown random model error. The assimilated observations are assumed to be related to the truth as
${\bf x}^{\rm obs}_{t}={\bf H}_{t}{\bf x}_{t}+\boldsymbol{\varepsilon}_{t},$ (2)
where ${\bf H}_{t}$ is the known observation-model operator and $\boldsymbol{\varepsilon}_{t}$ is the unknown random observation error.
If the errors $\boldsymbol{\eta}_{t}$ and $\boldsymbol{\varepsilon}_{t}$ are all mutually uncorrelated and have known first and second moments, then the mean-square optimal estimate of ${\bf x}_{t}$ given ${\bf x}^{\rm obs}_{t},{\bf x}^{\rm obs}_{t-1},\dots$ is given by the Kalman filter (e.g. Jazwinski, 1970). In high dimensions, however, the Kalman filter is prohibitively expensive. Its Monte-Carlo version known as the Ensemble Kalman filter (EnKF) is an affordable approximate filter, (e.g. Asch et al., 2016).
EnKF propagates in time a sample (ensemble) of ‘possible truths’ consistent with observations. A forecast (prior) ensemble is computed in EnKF at every assimilation cycle $t$ before assimilation of current observations, and an analysis (posterior) ensemble is computed after current observations are assimilated. The key assumption in EnKF is that the prior ensemble is a sample from the true prior distribution. EnKF computes a regularized sample covariance matrix $\widehat{\bf B}_{t}$ from the prior ensemble and substitutes it for the true prior covariance matrix ${\bf B}_{t}$ in the Kalman-filter analysis equation. It is the prior sample covariance matrix that we aim to regularize in this research using a spatial model — for systems whose state is a spatial field.
Having a deterministic forecast (the mean of the prior ensemble or the so-called control forecast), we deal in what follows with the the forecast-error random field, that is, the deviation of the forecast from the truth. Online information on its current (i.e., at the time instant $t$) probability distribution is given by ensemble perturbations, the differences of the prior ensemble members and the deterministic forecast.
For more details on approximate sequential filtering, see A.
@@ 3 Spatial model
We employ a non-parametric spatial model for the forecast-error random field to regularize the data assimilation problem. Following Higdon et al. (1999), we rely on the Gaussian process convolution model. In contrast to most applications, which postulate a parametric model for the spatial convolution kernel (e.g. Lemos and Sansó, 2009, Sham Bhat et al., 2012, Heaton et al., 2014, Li and Zhu, 2020), we choose a non-parametric approach to allow for variable shapes of spatial covariances. Non-parametric convolution kernels were considered by Tobar et al. (2015) and Bruinsma et al. (2022) in the stationary time series context. Here we propose a non-parametric non-stationary spatial convolution model on compact manifolds such as the sphere and the circle.
The default domain is the two-dimensional unit sphere ${\mathbb{S}}^{2}$ (the domain of interest in geosciences). We also formulate and test the model on the unit circle ${\mathbb{S}}^{1}$ (used as a lower-dimensional counterpart of the sphere). We refer to ${\mathbb{S}}^{2}$ and ${\mathbb{S}}^{1}$ just as the sphere and the circle in the sequel. We denote points on the sphere as $x,y,\dots$ and area elements simply by ${\rm d}x,{\rm d}y,\dots$ On the sphere, we use the terms stationarity and isotropy interchangeably. The terms field and process are also used interchangeably. We confine ourselves to band-limited processes because in the context of sequential filtering, the spatial field to be estimated inevitably has limited spatial resolution. We also assume that probability distributions we deal with in this study have densities with respect to the Lebesgue measure.
The rest of this section is structured as follows. We start by defining our Gaussian process convolution model, including the introduction of so-called local spectra. Then, we realize that to be identifiable from the prior ensemble, the model needs to be constrained. We propose the following four constraints: (i) local isotropy of the convolution kernel, (ii) positive definiteness of the non-stationary convolution kernel as a function of distance, (iii) local stationarity, (iv) smoothness of the local spectra.
@@ 3.1 General process convolution model
The (non-stationary in general) covariance function of the process $\xi(x)$ defined by Eq. (3) is $B(x,x^{\prime})=\operatorname{\mathbb{E}}\xi(x)\xi(x^{\prime})=\int_{D}w(x,y)w(x^{\prime},y)\,{\rm d}y$.
@@ 3.2 Constraining the model: strategy
In the formulation Eq. (3), the model is not identifiable, that is, the function $w(x,y)$ is not unique given the process covariance function $B(x,x^{\prime})$. Indeed, consider an isotropic convolution kernel $\psi(\rho(y,y^{\prime}))$, where $\rho(y,y^{\prime})$ stands for the great-circle distance between the points $y$ and $y^{\prime}$. Let the Fourier-Legendre transform of $\psi(\rho)$ be equal to one in modulus (see Eq. (6) for the definition of the Fourier-Legendre transform we use). Then, it is straightforward to see that the covariance function remains unchanged if $w(x,y)$ is convolved with $\psi$333 By the convolution of $w(\cdot,\cdot)$ with $\psi(\rho(\cdot,\cdot))$ we mean $w_{*}(x,y)=\int w(x,z)\,\psi(\rho(z,y))\,{\rm d}z$. . (The convolution with such a kernel is analogous to the multiplication by an orthogonal matrix in the space-discrete case.)
@@ 3.3 Locally isotropic kernel
$\xi(x)=\int_{D}u(x,\rho(x,y))\,\alpha(y)\,{\rm d}y.$ (5)
In the remainder of this subsection, we develop spectral representations of the process defined by Eq. (5) and its covariance function. We perform the spectral (Fourier-Legendre) expansion of $u(x,\rho)$ with $x$ being fixed:
@@
where $\ell_{\rm max}$ is the bandwidth, $\sigma_{\ell}(x)$ are real-valued functions (we call them spectral functions) and $P_{\ell}$ are the Legendre polynomials. We substitute $\rho=\rho(x,y)$ into Eq. (6) and apply the addition theorem of spherical harmonics,
where $Y_{\ell m}$ is the spherical harmonic, $\ell$ its degree (sometimes called the total wavenumber), $m$ its order (the meridional wavenumber), and the asterisk denotes complex conjugation.
As a result, Eq. (6) can be rewritten as
@@
Then, we make use of the spectral expansion of the white noise,
$\alpha(y)=\sum^{\ell_{\rm max}}_{\ell=0}\sum^{\ell}_{m=-\ell}\widetilde{\alpha}_{\ell m}\,Y_{\ell m}(y),$ (9)
where $\widetilde{\alpha}_{\ell m}$ are mutually uncorrelated complex-valued random variables with $\operatorname{\mathbb{E}}\widetilde{\alpha}_{\ell m}=0$ and ${\mathsf{Var}}\,\widetilde{\alpha}_{\ell m}=1$. More specifically, $\widetilde{\alpha}_{l0}$ are real-valued and all the other $\widetilde{\alpha}_{\ell m}$ are complex circularly symmetric random variables (e.g. Gallager, 2013, section 3.7) such that $\widetilde{\alpha}_{l,-m}=\widetilde{\alpha}_{\ell m}^{*}$.
Substituting Eqs. (8) and (9) into (5) and utilizing orthonormality of spherical harmonics, we obtain the basic spectral representation of the random process that satisfies Eq. (5):
@@
The spatial covariances $B(x,x^{\prime})=\operatorname{\mathbb{E}}\xi(x)\xi(x^{\prime})$ can readily be obtained from Eq. (10) by taking into account that all $\widetilde{\alpha}_{\ell m}$ are mutually uncorrelated and again applying the addition theorem of spherical harmonics. Equivalently, we may use Eq. (5) for this purpose so that
$B(x,x^{\prime})=\int_{D}u(x,\rho(x,y))\,u(x^{\prime},\rho(x^{\prime},y))\,{\rm d}y=\sum^{\ell_{\rm max}}_{\ell=0}\frac{2\ell+1}{4\pi}\,\sigma_{\ell}(x)\,\sigma_{\ell}(x^{\prime})\,P_{\ell}(\cos\rho(x,x^{\prime})).$ (11)
If $u(x,\rho)=u(\rho)$ or, equivalently, if all spectral functions $\sigma_{\ell}(x)$ do not depend on $x$, then $\xi(x)$ becomes stationary, in which case the process variance ${\mathsf{Var}}\,\xi(x)$ can be obtained by summing $\frac{2\ell+1}{4\pi}\sigma_{\ell}^{2}$ known as the variance spectrum (energy per degree $\ell$). Besides, $f_{\ell}=\sigma_{\ell}^{2}$ is often called the modal spectrum (energy per ‘mode’, i.e., per pair $(\ell,m)$). In the non-stationary case, in which $\sigma_{\ell}$ depends on $x$, we call $f_{\ell}(x)=\sigma_{\ell}^{2}(x)$ the local spectrum. Equation (11) shows that the non-stationary process variance can be computed in the same way as the stationary one: ${\mathsf{Var}}\,\xi(x)=\sum_{\ell}\frac{2\ell+1}{4\pi}\sigma_{\ell}^{2}(x)\equiv\sum_{\ell}\frac{2\ell+1}{4\pi}f_{\ell}(x)$.
Non-negative definiteness (i.e., validity) of the model’s non-stationary covariance function $B(x,x^{\prime})$ is established in B.
The formulation of the spatial model Eq. (10) on the circle is outlined in C.